Suppose you know that a company’s stock currently sells for $66.00 per share and the required return on the stock is 10 percent. You also know that the total return on the stock is evenly divided between capital gains yield and dividend yield.

Required:

If it’s the company’s policy to always maintain a constant growth rate in its dividends, what is the current dividend per share? (Do not round intermediate calculations. Round your answer to 2 decimal places (e.g., 32.16).)
Answer #1

dividend yield=Capital gains yield=(10/2)=5%

dividend yield=Dividend for next period/Current price

Dividend for next period=(66*5%)=$3.3

Hence current dividend=Dividend for next period*Present value of discounting factor(rate%,time period)

=$3.3/1.05

=$3.14(Approx).